A car was to be sold at a 20% discount, which amounted to $2,200. How much would the car sell for after the discount?
Answer by rapaljer(4551)

Are you saying that the discount amounted to $2,200? Assuming that this is what you meant,

Let x = original price of the car.

The discount is: 20% of x = 2200
.2x = 2200

Divide both sides of the equation by .2 in order to solve for x.

x = 2200/.2
x = 22000/2 = $11000
Now, x is the original price, so the selling price after the discount, subtract off the $2200, which makes the discount price $11000-2200 = $8800.

A faster way might be to say that after a 20% discount, you have to pay 80%, which is 4 times the discount. So multiply the discount $2200 times 4 = $8800.